Billit Corporation is considering leasing a machine to a client for five years. The machine sells for $100,000 cash. The incremental borrowing rate is 8%. The implicit rate is unknown. Payments are quarterly and start at the commencement of the lease. What is the quarterly lease payment? Show your work.

Quarterly

Calculation of Initial Lease Liability:

Quaterly Payment

Non-Lease Component*

Minimum Lease Payment

Present Value Factor

Present Value of Mininum Lease Payment
Answer #1

PV = -100,000

Rate=8/4=2%

N=5*4=20

PV of annuity due = Annuity + Annuity*(1-1/(1+rate)^(number of terms-1))/rate

100000= A+ A*(1-1/1.02^19)/0.02

100000=A + A*15.678462

100000 =16.678462 A

A= 100000/16.678462

A = $5995.76

Hence the quarterly payment is $5995.76